OpenStudy (anonymous):

The perimeter of a rectangle is 208 inches . The length exceeds the width by 28 inches . Find the length and the width

OpenStudy (anonymous):

\[l=w+28\] and \[2l+2w=208\] so \[l+w=104\] replace l by w+28 get \[w+28+w=108\] \[2w+28=104\] \[2w=76\] \[w=38\]

OpenStudy (anonymous):

width is 38, length is 38+28=66

OpenStudy (anonymous):

how did you get 104?

OpenStudy (anonymous):

oh i had \[2l+2w=208\] because the perimeter is twice the length plus twice the width. then to make life simpler i divided everything by 2

OpenStudy (anonymous):

oh Okay thank you !!!

OpenStudy (anonymous):

yw